SYLLABUS: MATH 788F

THE THEORY OF IRREDUCIBLE POLYNOMIALS

Instructor: Michael Filaseta

Office: 301 LeConte

Office Phone Number: 777-6589

Office Hours: By Appointment

Text Book: None

Class Notes: Notes will be made available at:

    http://www.math.sc.edu/~filaseta/gradcourses/Math788F/Math788F.html

Grading: Homework (50%), 1 Test (20%), Final (30%)

Attendance Policy: It is not intended that you take advantage of the fact that class notes are being made available and miss lectures. You are expected to attend each lecture. You will be considered absent if you miss more than ten minutes of a lecture. If you are absent for more than four lectures, your letter grade at the end of the course will be one-half of a letter grade lower than the grade determined from the homework, test, and final exam. If you are absent for more than eight lectures, you should withdraw from the course; if you do not, your letter grade at the end of the course will be one full letter grade lower than the grade determined from the homework, test, and final exam.

Homework Policy: Homework assignments must be turned in on time. Late homework will not be accepted. Although you are encouraged to work on homework together, you are expected to write up your homework solutions independently and without the aid of other people’s written solutions to the problems.